OpenStudy (anonymous):

The lengths of two sides of a right triangle are given. Find the length of the third side. Round to the nearest tenth if necessary. legs: 18 in. and 16 in.

OpenStudy (anonymous):

does it say which two sides? not the hypotenuse correct?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

the formula to find the sides of a right triangle is a^2+b^2=c^2 where a and b are the legs and c is the hypotenuse. so plug those in: 18^2+16^2=c^2 324+256=c^2 580=c^2 c=24.08

OpenStudy (anonymous):

24.1 to the nearest tenth.

OpenStudy (anonymous):

i ♥ your picture courtskye!